OpenStudy (anonymous):

What is the equation of the line that passes through the points (2, -4) and (0, 8)? y= 6x+ 8 y= 2x+ 8 y= -6x+ 8 y= -2x+ 8

OpenStudy (anonymous):

First you want to find the slope of the line between the two points. You can do this by finding the change in x and y between the two points.

OpenStudy (anonymous):

Can you tell me what the changes in x and y are?

OpenStudy (anonymous):

i think it y= 2x+ 8

OpenStudy (anonymous):

Or you could just plug in the points xD Yeah that sounds right.

OpenStudy (anonymous):

Oh nvm, it isn't that

OpenStudy (anonymous):

The slope is -6 because y changes by +12, while x changes by -2. (12/-2)=-6

OpenStudy (anonymous):

why the slope is -2 ?

OpenStudy (anonymous):

@Cha1234

OpenStudy (anonymous):

It shouldn't be -2, it should be -6

OpenStudy (anonymous):

AAAH! i'm confused

OpenStudy (anonymous):

Slope is equal to the change in y over the change in x.

OpenStudy (anonymous):

Y changes by +12, while x changes by -2

OpenStudy (anonymous):

That gives us 12/-2=-6

OpenStudy (nurali):

\[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] x1=2 x2=0 y1=-4 y2=8 \[m=\frac{ 8+4 }{ 0-2 }\] \[m=\frac{ 12 }{ -2 }\] \[m=-6\] \[(y-y _{1})=m(x-x _{1})\] \[y+4=-6(x-2)\] \[y+4=-6x+12\] \[y=-6x+ 8\]

OpenStudy (anonymous):

who's right?

OpenStudy (anonymous):

I'm telling you it's m=-6...